BRICS Notes Series, 1999

March 4, 2004

This document is also available as PostScript and DVI.

References

[NS-99-3] Abstract, PostScript, PDF.

Peter D. Mosses and David A. Watt, editors. Proceedings of the Second International Workshop on Action Semantics, AS ’99, (Amsterdam, The Netherlands, March 21, 1999), May 1999. iv+172 pp.

[NS-99-2] PostScript, PDF.

Hans Httel, Josva Kleist, Uwe Nestmann, and Antnio Ravara, editors. Proceedings of the Workshop on Semantics of Objects As Processes, SOAP ’99, (Lisbon, Portugal, June 15, 1999), May 1999. iv+64 pp.

[NS-99-1] Olivier Danvy, editor. ACM SIGPLAN Workshop on Partial Evaluation and Semantics-Based Program Manipulation, PEPM ’99, (San Antonio, Texas, USA, January 22–23, 1999), January 1999.